OpenStudy (anonymous):

the sum of 2 numbers is 1000 and the difference is 998. what are the 2 numbers

OpenStudy (anonymous):

If we let one of the numbers be 'x' and another be 'y' do you think you can write the two pieces of information as algebraic expressions?

OpenStudy (anonymous):

no...buts the question is an algebraic one...

OpenStudy (anonymous):

Okay, I'll help you with the first one. The first statement says that the two numbers add up to 1000. We can write this as: x + y = 1000 Where x and y are the two numbers. Do you think you can do the difference one?

OpenStudy (anonymous):

yes...then the second statement says x-y=998

OpenStudy (anonymous):

now what are x & y

OpenStudy (anonymous):

There are a few ways to do this. This site has some information on solving these kind of equations. http://www.allaboutcircuits.com/vol_5/chpt_4/11.html I will use the substitution method. (1) x + y = 1000 (2) x - y = 998 Rearrange 2 to get 'y' as the subject. x - y = 998 y = -998 + x Now substitute this new 'y' into the original (1). x + (-998 + x) = 1000 x -998 + x = 1000 2x = 1998 x = 999 Now, to find 'y', substitute this new x either (1) or (2). x - y = 998 999 - y = 998 y = 999 - 998 y = 1 Therefore the two numbers are 999 and 1.

OpenStudy (anonymous):

Does that help? :)

OpenStudy (anonymous):

super...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welome